OpenStudy (anonymous):

Carrie and crystal live at equal distances from school. Carrie walks to school at an average rate of 3 mph and crystal rides her bicycle at an averagte rate of 9 mph. it takes carrie 20 minutes longer than crystal to get to school. how far from school do crystal and carrie live? show work

OpenStudy (anonymous):

write let statement

OpenStudy (saifoo.khan):

y = distance from school (same for both) in miles x = how long it takes crystal to get to school in hours x+(1/3) = how long it takes carrie to get to school in hours Carrie: y miles = x+(1/3) hours * 3 mph Crystal: y miles = x hours * 9 mph Since both =y, then x+(1/3) hours * 3 mph=x hours * 9 mph x+(1/3) * 3 = x * 9 3x+1 = 9x 1 = 6x x = 1/6 hours = 10 minutes <---how long it takes crystal x + 1/3 = 1/6 + 1/3 = 1/6 + 2/6 = 3/6 = 1/2 hours = 30 minutes <---how long it takes Carrie